A football is kicked with a velocity of 20 m/s at an angle of 37° with the horizontal. Determine:


a. Maximum height reached by the ball


b. Time of flight


c. Horizontal distance traveled

1
Expert's answer
2021-11-22T16:16:04-0500

(a) Let's first find the time that the football takes to reach the maximum height:


"v=v_0sin\theta-gt,""0=v_0sin\theta-gt,""t=\dfrac{v_0sin\theta}{g}."

Then, we can find the maximum height reached by the ball from the kinematic equation:


"y_{max}=v_0tsin\theta-\dfrac{1}{2}gt^2."

Substituting "t" into the previous equation, we get:


"y_{max}=v_0sin\theta(\dfrac{v_0sin\theta}{g})-\dfrac{1}{2}g(\dfrac{v_0sin\theta}{g})^2,""y_{max}=\dfrac{v_0^2sin^2\theta}{2g},""y_{max}=\dfrac{(20\ \dfrac{m}{s})^2sin^237^{\circ}}{2\times9.8\ \dfrac{m}{s^2}}=7.39\ m."

(b) We can find the time of flight as follows:


"t_{flight}=2t=\dfrac{2v_0sin\theta}{g},""t_{flight}=\dfrac{2\times20\ \dfrac{m}{s}\times sin37^{\circ}}{9.8\ \dfrac{m}{s^2}}=2.46\ s."

(c) We can find the horizontal distance traveled as follows:


"x=v_0t_{flight}cos\theta=20\ \dfrac{m}{s}\times2.46\ s\times cos37^{\circ}=39.3\ m."